Use the density formula to solve the following problem:
A student finds a rock on the way to school. In the laboratory he determines that the volume of the rock is 29.7 cm3, and the mass in 139.943 g. What is the density of the rock?
1
Expert's answer
2017-02-21T09:41:42-0500
D = M / V (D = Density, M = Mass, V = Volume)
V=29.7 cm3 and M=139.943g
Density of the rock:
D=139.943g/29.7cm3=4.71 g/cm3

Answer: 4.71 g/cm3